please help i do not get this exercise problem thank you so much in advance
L has y-intercept (0, _3) and is parallel to the line with equation y= 2/3x+1

Answer by malakumar_kos@yahoo.com(315)


L has y-intercept (0, _3) and is parallel to the line with equation y= 2/3x+1
Let the eq'n to the line L is y = Mx+c
the line is parallel to y = 2x/3+1 therefore their slopes are equal
the slope of th line L is M = 2/3
to find th value of c substitute the value of x,y & M in y = Mx+c
-3 = (2/3).0+c
-3 = 0+c or c= -3
The required eq'n is y = 2/3x-3
3y = 2x-9 (multiply by 3 throughout)